# Spurious reconstruction from brain activity

### v1

@@ 2.2.1 Output dimension collapse
Next, we consider a scenario where the diversity of the target features is small. This situation can arise when the feature space exhibits a clustered structure and the training data lacks sufficient diversity, as observed in the case study with the CLIP text features and the NSD dataset. When the training feature values have limited diversity, the predicted values from brain activity, which are represented as linear combinations of these features, also become constrained. Consequently, the prediction from brain data to features effectively becomes a projection onto a low-dimensional subspace formed by the training data.
To illustrate this phenomenon, we conducted a simple simulation to examine the distribution of predicted values from a linear regression model trained on clustered features (Figure 8). We generated clustered features by sampling from a Gaussian mixture distributions in a high-dimensional space. The corresponding brain activity data was generated from a multivariate Gaussian distribution. We then trained a linear regression model to predict the clustered feature values from the randomly generated brain data.
Figure 8: Output dimension collapse. The left panel shows the data distribution of source brain activity in the first two dimensions of high-dimensional space. The right panel shows the distribution of target latent features in the first two principal component (PC) dimensions. If the target features form a cluster structure in the high-dimensional space (left panel green points), the predictions (gray colors) fall into a subspace created by the training clusters, regardless of the input data (right panel).
@@
The simulation results clearly demonstrate the impact of clustered features on the predicted values. The trained model projects arbitrary brain data onto the subspace defined by the training features, resulting in predicted values that are confined to the vicinity of the training clusters. This observation highlights the limitation of training linear regression models with clustered features: the trained model’s predictions are inherently constrained by the diversity and structure of the training data.
@@
This phenomenon, which we refer to as “output dimension collapse,” has important implications for the generalization capability of linear regression models in the context of brain decoding and visual image reconstruction. When the training data lacks diversity and forms distinct clusters in the feature space, the translator overly adapts to the subspace formed by the training data, regardless of the potential of the latent feature space. Consequently, the translator’s outputs become similar to patterns in the training set, irrespective of the inputs, severely limiting the model’s ability to predict novel or out-of-distribution samples. Note that output dimension collapse is not inherently caused by the limitations of linear regression models but rather stems from a lack of diversity in the outputs of the training data. This phenomenon can occur regardless of the type of regression model employed, including nonlinear models. In fact, nonlinear models may be even more susceptible to output dimension collapse, as the increased flexibility and complexity of these models can potentially exacerbate the issue.
@@
Output dimension collapse may explain why plausible reconstructions were obtained even from random brain data by merely generating images several times in the case study. The NSD’s lack of semantic diversity causes the translator to adapt only to the feature patterns of the training set, restricting its outputs to the subspace formed by the training data. As a result, convincing images could be found even from random data through questionable post-hoc selection.
@@
2.2.2 Simulation with clustered features: What makes prediction compositional
@@
The case study revealed that the NSD exhibits limited diversity (Figure 4) and poses difficulties for zero-shot prediction (Figure 5 and 6). These observations suggest that the translator of CLIP features suffers from output dimension collapse due to the lack of semantic diversity in NSD. To explore potential strategies for mitigating output dimension collapse and achieving flexible predictions, we conducted simulation analyses using clustered data to assess generalization performance beyond the training set.
@@
Our simulation involved teacher-student learning with Gaussian mixture models (see Materials and Methods: “Simulation with clustered data”). To imitate the feature translation situation from brain activity, we first generated feature data and assumed a scenario where observation noise affects the brain data. To simulate a situation where the dataset has cluster structures and to control diversity effectively, the training feature $\mathbf{y}\in\mathbb{R}^{D}$ was generated from a $D$-dimensional mixture of Gaussians (Figure 9a). The brain activity data $\mathbf{x}\in\mathbb{R}^{D}$ was generated from $\mathbf{y}$ multiplied by teacher weight and adding observation noise.
Figure 9: Simulation analysis of predicting features with cluster structure. (a) Illustration of target latent features setting. The latent features were generated from Gaussian mixture distributions. $\sigma_{r}^{2}$ is a parameter of the sample variance of each cluster. $\sigma_{s}^{2}$ controls the distance across clusters. In-distribution test samples are generated from the same clusters as in the training set. Out-of-distribution (OOD) test samples are generated from a novel cluster not included in the training set. (b) Cluster identification accuracy. The $x$ axis represents the number of clusters used in the training and the $y$ axis represents the cluster identification performance. Green line shows the results of In-distribution test samples and orange line shows those for OOD samples. Dashed curve indicates the chance level. (c) Illustration of a possible scenario where sufficient data diversity is needed to achieve generalizability. The upper illustration shows that the training data should cover the whole latent feature space, requiring an exponential order relative to the feature dimension. The lower one shows that the training data is covering only the effective axes of the latent feature space, leading to a linear order relative to the feature dimension. (d) Relation between latent feature dimension and sufficient number of clusters for generalization. The $x$ axis represents the dimension of target latent features. The $y$ axis represents the number of clusters that achieve above $0.5$ cluster identification accuracy. Three lines indicate the results from different $\sigma_{s}^{2}/\sigma_{r}^{2}$. (e) Representative example of generalizability beyond the training set. Light green points indicate the in-distribution test samples while light orange points indicate OOD test samples. Green and orange points represent the predictions.
@@
We trained a ridge regression model on large training data samples and obtained the student weight. To simulate the situation where the trained model encounters clusters that are not available at the training phase, we generated two types of test samples, in-distribution and out-of-distribution. In-distribution test samples were generated using the same probability distribution as the training data, while out-of-distribution (OOD) samples were generated from a different distribution. For these two types of predicted features, we calculated the cluster identification accuracy (Figure 6c), using $C+1$ cluster centers, which included $C$ centers from the training set and one additional center for the OOD test set.
@@
We first examined zero-shot prediction performance as the number of clusters in the training data was increased while fixing the feature dimension $D$ and the cluster variance ratio between each cluster constant (Figure 9b). While the cluster identification performance of in-distribution samples were perfect, that of OOD samples showed different patterns depending on the number of clusters in the training data. When the number of training clusters was low, cluster identification accuracy was $0\%$. This indicates that the behavior of the translator became more similar to a classifier, difficult to generalize beyond the training set, as observed in NSD cases (see also Figure 6d). On the other hand, as the number of clusters in the training data increased, it became possible to identify the novel clusters, achieving the same performance as in-distribution test samples. This observation indicates the importance of the diversity of the training dataset. We also emphasize that large numbers of training data do not necessarily address the problem. All of these results were obtained with a sufficiently large number of training data, and the number of training clusters was varied while keeping the number of training data fixed. Also, we observed qualitatively similar results in increasing the data diversity by controlling the cluster variance ratio (Supplementary Figure S3a).
@@
Next, we investigate how diverse the training data needs to be to ensure sufficient generalization. There are two possible scenarios for diversifying training samples: either by densely sampling the entire space so that there are no remaining gaps or by uniformly sampling to the extent that it covers the entire dimension of the feature space (Figure 9c). The former scenario requires an exponentially larger number of samples relative to the dimension while the latter scenario only requires up to a linear order. We sought to reveal which of these two scenarios was more likely to be true by varying the dimensions of the feature space and identifying the number of clusters required for generalization (see also Supplementary Figure S3b). Here, we defined the number of clusters required for generalization as the point at which the prediction accuracy of OOD samples exceeds $50\%$. The relationship between the dimension of the feature space and the number of clusters necessary for generalization appears to be linear (Figure 9d). This finding suggests that achieving generalization does not necessarily require an exponentially large diversity that fills the entire feature space. Instead, it suffices to have a number of clusters that cover the effective dimensions within the feature space. Although obtaining a large amount of brain data is hard work, it is important for a dataset to contain sufficiently diverse stimuli covering the effective dimensions of the target feature space to achieve zero-shot prediction.
@@
We also confirmed this phenomenon by a simple and transparent example ($D=2$, Figure 9e). The training data covers sufficient axes in the feature space, enabling the prediction of locations not present in the training set. Based on this low dimensional intuition, we argue that successful zero-shot prediction requires that training data leads to representations that can serve as a basis spanning the feature space. Leveraging such bases effectively enables the model to predict novel samples by predicting each basis and combining them. This compositional representation, spanning the target feature space is crucial for zero-shot prediction (Schug et al., 2024) and reconstructing arbitrary visual images from limited brain data.
2.2.3 Preserved image information across hierarchical DNN layers
To enable the reconstruction for arbitrary visual images, compositional latent features that can be appropriately mapped into the image space are essential. Although it is often assumed that the hierarchical feature representations in DNNs, especially convolutional neural network (CNN), discard pixel-level information through their hierarchical processing with progressively expanding receptive fields, this is not necessarily the case.
For example, the latent features of auto-encoder models (Hinton and Salakhutdinov, 2006; Kingma and Welling, 2014; van den Oord et al., 2017) can represent images in a low-dimensional space while preserving their reversibility, which is reasonable considering that the output is trained to match the input. Furthermore, Mahendran and Vedaldi (2015) showed that input images can be reconstructed with reasonable accuracy even from relatively high-level layers of a DNN designed for an object recognition task. It has also been argued that large receptive field sizes do not necessarily impair neural coding capacity as long as the number and density of units remain constant (Zhang and Sejnowski, 1999; Majima et al., 2017). These results challenge the notion that higher-level layers in DNNs discard all pixel-level information and highlight the potential for utilizing intermediate representations in visual image reconstruction. To further illustrate this point, we performed a recovery check on each intermediate layer of the VGG19 network used in the iCNN methods (Figure 10; see also Figure 7). Given the DNN features of a target image, we optimized input pixel values to make the image’s latent features similar to the targets (See Materials and Methods “Recovery check of a single layer by iCNN”). We observed that input images can be recovered with reasonable accuracy from relatively high-level layers (around the 11th layers, the 19th layers as a total). Furthermore, by introducing image generator networks to add constraints on image statistics, reasonable recovery can be achieved from even higher layers. By utilizing a weak image prior (Ulyanov et al., 2017), which contains only information about the structure of images without any prior information on natural images, input images can be recovered from the 13th layer. When using an image generator that has learned natural image information (Shen et al., 2019b; Dosovitskiy and Brox, 2016), input images can be recovered even from the 15th layer.
Figure 10: Recovery check: single layer used in iCNN method. For three images, the different optimization algorithm is applied to recover the visual input. The left one is pixel optimization, which directly optimizes the pixel values so that the loss between that generated image feature and target image features is minimized. The middle one is optimization with week image prior, which optimizes the parameters of image prior models and their latent features simultaneously (see Material and methods). The right one is optimization with pretrained image prior, which optimizes the latent features of the fixed image generator model so that the loss between the output image features and target image features is minimized.
These findings underscore the significance of intermediate DNN representations that can effectively recover the input image in visual image reconstruction studies. Conversely, utilizing high-level image features or features from other modalities, such as text annotations, for which the corresponding images are challenging to recover, is not a rational choice for visual image reconstruction tasks. Although recent text-to-image models and decoded text features can easily generate images, the outputs should not be interpreted as reconstruction results. Instead, it is more reasonable to interpret them as visualizations of decoded semantic information.
2.2.4 Caveat with evaluation by identification
Figure 11: Pairwise sample identification with structured data. The $x$ axis represents the number of categories in the test set. The $y$ axis represents the pairwise sample identification accuracy. The lines show the different accuracy $q$ in the situations where the candidate sample is from a different category than those in the test samples. The representative similarity matrices corresponding to these conditions are shown in the figure above.
@@ 2.2.4 Caveat with evaluation by identification
Pairwise identification has been a standard metric for evaluating latent feature decoding (Horikawa and Kamitani, 2017) or reconstruction performance (Beliy et al., 2019; Shen et al., 2019a, b; Mozafari et al., 2020; Qiao et al., 2020; Ren et al., 2021; Gaziv et al., 2022; Takagi and Nishimoto, 2023b; Ozcelik and VanRullen, 2023; Ferrante et al., 2023; Scotti et al., 2023; Denk et al., 2023; Koide-Majima et al., 2024). However, our analysis revealed that even with difficulties in accurately identifying semantic clusters the test samples belong to, pairwise identification performance still surpassed chance levels (Figure 6d). Here we critically examine this metric and demonstrate that significant results can be easily obtained when the target or predicted features exhibit certain structures.
@@
Pairwise identification is calculated as the accuracy with which the predicted features (reconstructed image or its features) can correctly identify the true ones, in pairs consisting of a true sample and one of the remaining samples in the test set. If the candidate feature belongs to the same category as the true feature, we can expect the identification to be difficult. Conversely, if the candidate feature belongs to a different category from the true feature, identification would become easier.
@@
Here we assume the test set comprises $k$ categories and all test samples are equally distributed across each category for simplicity. We model the situation mentioned above as the expected identification accuracy is set to chance level (i.e., 0.5) when a candidate belongs to the same category as the true feature. In contrast, the expected identification accuracy when a candidate belongs to a different category from the true feature is set to $q$, ranging between 0.5 to 1 to mimic the ease of identification across categories. Assuming the number of test samples is large enough, the pairwise identification becomes

### v3

@@ Results
Next, we consider a scenario where the diversity of the target features is small. This situation can arise when the feature space exhibits a clustered structure and the training data lacks sufficient diversity, as observed in the case study with the CLIP text features and the NSD dataset (Fig 4A). When the training features have limited diversity, the predicted values from brain activity, which are represented as linear combinations of these target features, also become constrained. Consequently, the prediction from brain data to target features effectively becomes a projection onto a low-dimensional subspace formed by the training data.
Fig 8: Demonstration of output dimension collapse in feature prediction. The left panel shows the distribution of source brain activity data in the first two dimensions of the original high-dimensional space. The right panel displays the distribution of target latent features projected onto the first two principal component (PC) dimensions. A linear ridge regression model is trained using the training data (gray points in the left and right panels), in which the output features in the training data are constrained to two clusters. When presented with test data from the brain space (left panel), the model generates predictions (orange points) in the latent feature space (right panel). This visualization demonstrates how the predicted features are constrained to the subspace defined by the training data, highlighting the limitation in generalizing beyond the training data distribution.
To illustrate this phenomenon, we simulated teacher-student learning, a framework where a “teacher” model generates data based on certain underlying rules, and a “student” model is trained to learn or approximate those rules by observing the generated data. We examined the distribution of predicted values from the student linear regression model trained on clustered features generated by the teacher model. We generated clustered features by sampling from a Gaussian mixture distribution in a high-dimensional space. The corresponding brain activity samples were generated from the latent feature samples multiplied by teacher weight and adding observation noise (See Materials and Methods “Simulation with clustered data”). Student weights were obtained by training a linear regression model to predict the clustered feature values from the corresponding brain data. We then projected the randomly generated brain samples into latent feature space by learned regression weight. To visualize the high-dimensional predicted patterns effectively, we projected the predicted features onto the PCA spaces derived from the training features and showed the first two dimensions.
@@
The simulation results clearly demonstrate the impact of clustered features on the predicted values (Fig 8). The trained linear regression model projects arbitrary brain data onto the subspace defined by the latent features in the training set, resulting in predicted values that are confined to the vicinity of the training clusters. This observation highlights the limitation of training linear regression models with clustered features as prediction targets: the trained model’s predictions are inherently constrained by the diversity and structure of the training data.
@@
We refer to this phenomenon as “output dimension collapse,” where the model’s predictions become confined to a limited subspace of the output feature space. It has important implications for the generalization capability of linear regression models in the context of brain decoding and visual image reconstruction. When training data lack diversity and form distinct clusters in the feature space, the translator overly adapts to the subspace formed by the training data, regardless of the potential of the latent feature space. Consequently, the translator’s outputs become confined to patterns similar to those in the training set, severely limiting the model’s ability to predict novel or out-of-distribution samples.
@@
Output dimension collapse may explain why plausible reconstructions were obtained even from random brain data by merely generating images several times in the case study (Fig 3). The lack of semantic diversity in NSD causes the translator to adapt only to the feature patterns of the training set, restricting its outputs to the subspace formed by the training data. As a result, convincing images could be found even from random data through questionable post-hoc selection.
It is crucial to understand that output dimension collapse may not inherently be a limitation of linear regression models. This phenomenon can manifest across various regression models, including nonlinear approaches such as multi-layer perceptrons (MLPs) [46] and transformers [48]. Nonlinear models may be more susceptible to output dimension collapse due to their enhanced flexibility in adapting to training data.
It is also important to recognize that the mathematical formulation in Eqs 1–3 assumes that all input variables are shared across target variables. If each target is predicted from a distinct set of input variables through feature (voxel) selection, the predictions can become more independent, potentially mitigating output dimension collapse. This approach has been utilized in the field since its early days, with techniques such as sparse voxel selection and modular modeling [28, 29, 61, 37].
@@
Simulation with clustered features: What makes prediction compositional?
@@
The case study revealed that the NSD exhibits limited diversity (Fig 4) and poses difficulties for zero-shot prediction (Fig 5 and Fig 6D). These observations suggest that the translator of CLIP features suffers from output dimension collapse due to the lack of semantic diversity in the NSD. To explore potential strategies for mitigating output dimension collapse and achieving flexible predictions, we conducted simulation analyses using clustered features to assess generalization performance beyond the training set.
@@
As in the previous section, our simulation involved teacher-student learning. We first generated feature data ${\mathbf{y}}\in\mathbb{R}^{D}$ then made the input brain data ${\mathbf{x}}\in\mathbb{R}^{D}$ by translating ${\mathbf{y}}$ with the teacher weight and added observation noise. To simulate a situation where the dataset has cluster structures and to control diversity effectively, the training feature vector $\mathbf{y}\in\mathbb{R}^{D}$ was generated from a $D$-dimensional Gaussian mixture (Fig 9A).
@@
We trained a ridge regression model on large training data samples and obtained the student weight. To simulate the situation where the trained model encounters clusters that are not available at the training phase, we used two types of test samples, in-distribution and out-of-distribution test samples: in-distribution test samples were generated from one of the clusters used in the training set, whereas out-of-distribution (OOD) samples were generated from the novel cluster that is not included in the training set. For these two types of predicted features, we calculated the cluster identification accuracy (Fig 6C) by using $C+1$ cluster centers: $C$ centers from the training set and one cluster center of the OOD test set.
Fig 9: Simulation analysis of predicting features with a cluster structure (A) Illustration of target latent features setting. The latent features were generated from Gaussian mixture distributions. $\sigma_{\mathrm{intra}}$ is a parameter that controls the scale of the samples within a cluster. $\sigma_{\mathrm{inter}}$ controls the scale between the cluster centers. In-distribution test samples are generated from one of the clusters in the training set. Out-of-distribution (OOD) test samples are generated from a novel cluster not included in the training set. (B) Cluster identification accuracy. The $x$-axis represents the number of clusters used in the training, and the $y$-axis shows the cluster identification performance. The green line represents the results of in-distribution test samples, and the orange line shows the results for OOD samples. The dashed curve indicates the chance level. (C) Scenarios for achieving generalizability with sufficient data diversity. The upper illustration shows the training data covering the whole latent feature space, requiring an exponential order relative to the feature dimension. The lower one shows that the training data covers only the effective axes of the latent feature space, leading to a linear order relative to the feature dimension. (D) Relation between latent feature dimension and the sufficient number of clusters for generalization. The $x$-axis represents the dimension of the target latent features. The $y$-axis shows the number of clusters that achieve above $0.5$ cluster identification accuracy. The three lines indicate the results from different ratios of $\sigma_{\mathrm{intra}}^{2}/\sigma_{\mathrm{inter}}^{2}$. (E) Example of successful prediction beyond the training distribution in a 2D output feature space. Dark green points represent the in-distribution target features, with their corresponding light green points showing the model’s predictions. Dark orange points denote the OOD target features, with light orange points indicating their predictions. Despite the OOD target features (dark orange) not being included in the training clusters, they are accurately predicted (light orange), demonstrating the model’s generalization ability. This successful generalization highlights the power of compositional prediction, where the model effectively combines individual dimensions to predict novel feature combinations outside the training distribution.
@@
We first examined zero-shot prediction performance for different numbers of training clusters (Fig 9B). We fixed the feature dimension $D$ and the cluster variance ratio constant. While the cluster identification performances of in-distribution test samples were perfect, the performances of OOD test samples showed different patterns depending on the number of clusters in the training data. When the number of training clusters was small, the cluster identification accuracy was 00%. This result indicates that the behavior of the translator became more similar to a classifier, making it difficult to generalize beyond the training set, as observed in NSD cases (see also Fig 6D). On the other hand, as the number of clusters in the training data increased, it became possible to identify the novel clusters, achieving the same performance as in-distribution test samples. This observation indicates the importance of the diversity of the training dataset. We also emphasize that large numbers of training samples do not necessarily address the problem. All of these results were obtained with a sufficiently large amount of training data, and the number of training clusters was varied while keeping the amount of training data fixed. Also, we observed qualitatively similar results in increasing the data diversity by controlling the cluster variance ratio and keeping the number of dimension and training clusters constant (S4 Fig.A).
@@
Next, we investigate how diverse the training data needs to be to ensure sufficient generalization. There are two possible scenarios for diversifying training samples: either by densely sampling the entire target feature space so that there are no remaining gaps (Fig 9C; top) or by uniformly sampling to the extent that it covers the entire dimension of the target feature space (Fig 9C; bottom). The former scenario requires an exponentially larger number of samples/clusters relative to the dimension, whereas the latter only requires up to a linear order. We sought to reveal which of these two scenarios was more likely to be true by varying the dimensions of the feature space and identifying the number of clusters required for generalization (S4 Fig. B for identification accuracy in each condition). Here, we defined the number of clusters required for generalization as the point at which the identification accuracy of OOD samples exceeds $50$%. The relationship between the dimension of the feature space and the number of clusters necessary for generalization appears to be linear (Fig 9D). This finding suggests that achieving generalization does not necessarily require an exponentially large diversity that fills the entire feature space. Instead, it suffices to have a number of clusters that cover the adequate dimensions within the target feature space. Although obtaining a large amount of brain data is hard work, it is important for a dataset to contain sufficiently diverse stimuli covering the effective dimensions of the target feature space to achieve zero-shot prediction.
@@
We also confirmed this phenomenon with a simple and transparent example ($D=2$, Fig 9E). The training data covers sufficient axes in the target feature space, enabling the prediction of locations not present in the training set. Based on this low-dimensional intuition, we argue that successful zero-shot prediction requires that training data leads to representations that can serve as a basis for spanning the target feature space. Leveraging such bases effectively enables the model to predict novel samples by predicting each basis and combining them. This compositional representation, spanning the target feature space, is crucial for zero-shot prediction [62] and reconstructing arbitrary visual images from limited brain data.
Caveat with evaluation by pairwise identification
@@
Pairwise identification has been a standard metric for evaluating latent feature decoding [19] or reconstruction performance [39, 29, 38, 34, 36, 40, 41, 44, 45, 46, 57, 58]. However, our analysis revealed that even with difficulties in accurately identifying semantic clusters the test samples belong to, the pairwise identification performance still surpassed its chance level (Fig 6F). Here we critically examine this metric and demonstrate that significant results can be easily obtained when the target or predicted features exhibit certain structures.
@@
Pairwise identification is calculated as the accuracy with which the predicted features (either the output of a translator or features extracted from the output of a generator) can correctly identify the corresponding true ones, in pairs consisting of a true sample and one of the remaining samples in the test set. If the candidate sample belongs to the same category as the true sample, the identification is expected to be difficult. Conversely, if the candidate sample belongs to a different category from the true one, identification becomes easier.
Fig 10: Expected pairwise sample identification performance in categorically structured data. The $x$-axis represents the number of categories in the test set. The $y$-axis represents the pairwise sample identification accuracy. Different lines represent various levels of accuracy ($q$) in distinguishing samples from different categories. The chance level of $0.5$ is represented by the bottom line. Above the graph, hypothetical similarity matrices are shown to illustrate the categorical structures of the samples, where the color (white/black) indicates similarity between samples. The block diagonal structure reflects the categorical nature of the data. Samples within the same category are assumed to be indistinguishable, resulting in a pairwise identification accuracy of $0.5$ (chance level). Samples from different categories can be distinguished with a pairwise identification accuracy of $q$, where q varies between $0.5$ and $1$.
@@
In this model, we assume the test set comprises $k$ categories with test samples equally distributed across each category for simplicity. We model the situation mentioned above by setting the expected identification accuracy at the chance level (i.e., $0.5$) when a candidate sample belongs to the same category as the true one. Conversely, when the candidate sample belongs to a different category than the true sample, we set the expected identification accuracy to a parameter $q$, ranging from 0.5 to 1. This parameter $q$ reflects the ease of identification across categories. Assuming the number of test samples is sufficiently large, the pairwise identification accuracy $\mathrm{Acc}$ becomes